# Simulation of time crystal behavior for a few boson chiral soliton model in a ring

### v3

@@ Abstract
We present numerical simulations for a chiral soliton model with $N=2,3$ bosons in a ring, this being a few-particle version of our previous mean-field model for a quantum time crystal OW19 . Following Syrwid, Kosior, and Sacha (SKS), the notion is that a precise position measurement of one particle can lead to spontaneous formation of a bright soliton that in a time crystal should rotate intact for at least a few revolutions around the ring SKS_EPL . In their work SKS find spontaneous formation of a soliton due to the position measurement, but quantum fluctuations cause the soliton to subsequently decay before it has a chance to perform even one revolution of the ring. Based on this they conclude that time crystal dynamics are impossible. In contrast, for our few boson chiral soliton model, and allowing for imprecise (weak) measurements of the particle position, we show that time crystal behavior is possible allowing for several revolutions of the spontaneously formed soliton around the ring. We therefore argue that our chiral soliton model can realize a quantum time crystal when weak position measurements are allowed for.
@@ I Introduction
The issue undermining Wilczek’s initial proposal was nicely illustrated in a paper by Syrwid, Kosior, and Sacha (SKS) in which they simulated N-bosons in a ring that is pierced by a flux tube, and which was initiated from the N-particle ground state which shows no signatures of rotation of localization SKS_EPL . What they found was that precise position measurement of one particle could lead to the spontaneous formation of a $(N-1)$-particle soliton that is rotating, but that quantum fluctuations cause this soliton to subsequently decay before it has a chance to perform even one revolution of the ring. SKS presented results for particle numbers in the range $N=10-60$, and the lifetime $t_{c}$ of the soliton was found to decrease with $N$. We observed the same decay for a non-chiral version of our model with $N=3$, both for precise and imprecise position measurements (with an uncertainty in the position measurement), the lifetime being longer for the imprecise measurement. The conclusion of SKS remains intact that Wilczek’s model cannot show time crystal behavior SKS_EPL .
@@
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 . This paper was followed by discussion in the literature with SKS SKS_comment -OW_lack , culminating in their paper highlighted in the previous paragraph SKS_EPL . The goal of the present paper is to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time crystal behavior is possible. In particular, we find the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson soliton can persist over several revolutions of the ring as required for time crystal behavior. We find that for an imprecise or weak position measurement time crystal behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS SKS_EPL . We therefore argue that our chiral model can realize a quantum time crystal when weak position measurements are given due consideration.
@@
The remainder of this paper is organized as follows: Section II describes the few boson version of our chiral soliton model along with a description of the parameters involved. The case of $N=2$ particles is covered in Sec. III and shows the ideas involved in a transparent manner. Section IV contains the main results and covers the three particle case that demonstrates that our few boson chiral soliton model can display time crystal behavior in which a two-boson soliton state Jackiw can execute several revolutions of the ring. Summary and conclusions are given in Sec. V.
@@ II.1 N-particle Schrödinger equation
$i{\partial\Psi\over\partial t}=\sum_{j=1}^{N}\left(-i{\partial\over\partial\theta_{j}}-A(\theta_{1},\theta_{2},\ldots\theta_{N})\right)^{2}\Psi+{g\over 2}\rho(\theta_{1},\theta_{2},\ldots\theta_{N})\Psi.$ (1)
@@ II.2 Model parameters
In our previous mean-field model the chiral soliton model was derived starting from the spinor Schrödinger equation for a system of two-level bosonic atoms trapped in a ring that are dipole coupled using a laser beam carrying orbital angular momentum (OAM) characterized by the winding number $\ell$. By preparing the atoms in the appropriate dressed state a scalar field analysis may be obtained in which the atoms are subject to a vector potential of the form in Eq. (2) with $A^{(0)}={\ell\over 2}$. Moreover, the density-dependent gauge potential results from a collision-induced detuning, with the consequence that the dimensionless parameter $\kappa\propto\ell$. This means that if the winding number of the laser beam $\ell=0$, then $\kappa=0$ and there will be no chiral dynamics and no possibility for time crystal behavior. The term proportional to the dimensionless parameter $g$ describes energy-shifts due to many-body interactions.
@@ III.1 Two-particle Schrödinger equation
$i{\partial\Psi\over\partial t}=\sum_{j=1}^{2}\left(-i{\partial\over\partial\theta_{j}}-{\ell\over 2}-\kappa\eta(\theta_{1}-\theta_{2}))\right)^{2}\Psi+{g\over 2}\eta(\theta_{1}-\theta_{2})\Psi.$ (6)
@@
$i{\partial\Psi\over\partial t}=\left[-2{\partial^{2}\over\partial x^{2}}-{1\over 2}{\partial^{2}\over\partial s^{2}}+2i\left({\ell\over 2}+\kappa\eta(x)\right){\partial\over\partial s}+2\left({\ell\over 2}+\kappa\eta(x)\right)^{2}+{g\over 2}\eta(x)\right]\Psi.$ (7)
@@ III.2 Two-particle ground state
In the remainder of this section we choose $g=-2$ as a illustrative example, though similar results could be obtained for different combinations of parameters. Figure 2 shows some more detail of the solution for this choice. In particular, Fig. 2(a) shows the scaled energy versus $p$ and reveals that the ground state occurs for $p=\ell=2$, and (b) shows a line plot of the ground state profile $|\varphi(x)|^{2}$ (properly normalized) versus $x$ for $p=2$. Furthermore, the width (half-width at half-maximum) of the ground state profile in Fig. 2(b) with respect the the relative coordinate $x$ of the two particles is $w\sim 1$, the parameter $w$ characterizing the expected width of any localized two-particle solutions that may arise.
Figure 3: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2})|^{2}$, and (b) simulation for a measurement with $n=50$ that localizes particle $2$ with an uncertainty $\Delta\theta\sim 0.2$. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it does not reflect time crystal behavior. Parameters are the same as as in Fig. 1 along with $g=-2$.
Figure 4: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining particle after the measurement of particle 2 to a non-dispersing solution for the ring. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as as in Fig. 1 along with $g=-2$.
III.3 Time crystal dynamics from measurement
To proceed we first recognize that the ground state represented by the probability density in Fig. 2(b) does not correspond to a localized state on the ring since $|\varphi(x)|^{2}$ depends only on the relative coordinate $x$, and there is no information concerning the COM position $s$ of the particle. We can probe further and reconstruct the ground state two-body wave function $\Psi(\theta_{1},\theta_{2})$ (at $t=0$), where $|\Psi(\theta_{1},\theta_{2})|^{2}$ is shown in Fig. 3(a). From this it is straightforward to see that the probability density of finding the particle at $\theta_{1}$ for any value of $\theta_{2}$,
@@ III.3 Time crystal dynamics from measurement
where the function $G(\theta_{2}-\theta_{2}^{(0)})$ reflects the uncertainty of the position measurement of particle $2$, and is a distribution which obeys the periodicity of the ring, and with an adjustable width $\Delta\theta$. Here for illustration we employ the following distribution
@@
with ${\cal N}$ a normalization constant. In the limit $n\gg 1$ this distribution approaches a delta-function which corresponds to the precise measurement model adopted in the paper of SKS, whereas for $n=1$ the measurement uncertainty is around $\Delta\theta\sim 2$. After the measurement, precise or imprecise, the wave function $\psi(\theta_{1},t=0)$ (suitably normalized) for the remaining particle is given by Eq. (11), and subsequent time development is governed by the single-particle Schrödinger equation
@@
Figure 3(b) shows a simulation of the quantum dynamics according to Eq. (13) following a precise measurement for which we set $\theta_{2}^{(0)}=0$ and $n=50$ so that $\Delta\theta\sim 0.2$. In this case the uncertainty is much less than the expected width $w\sim 2$ for a two-particle localized state, so the measurement is expected to be very disruptive. The figure shows a color coded plot of the probability density $|\psi(\theta_{1},t)|^{2}$ versus $t$ and $\theta_{1}$. Then, for example, if we select $\theta_{1}=\theta_{2}^{(0)}=0$, and trace the probability density in time along the horizontal axis, we see what might be a periodic variation indicative of a time crystal. But this is an illusion based on us choosing $\theta_{1}=\theta_{2}^{(0)}$, a quantity we cannot know a priori, and if we chose another $\theta_{1}$ at which to monitor the time variation of the probability density we would generally find another much noisier variation with time. Another way to see this from Fig. 3(b) is that the spatial density profile $|\psi(\theta_{1},t=0)|^{2}$ becomes rapidly distorted as time increases, not simply displaced as expected for a time crystal. This is in perfect keeping with the findings of SKS that too precise a position measurement of one particle can produce a localized state initially but this will become rapidly dispersed before the initial localized solution can exhibit a single rotation around the ring.
@@
For time crystal behavior what is required is that the center-of-mass of the initial localized state translates in time while maintaining its shape for several transits of the ring. An estimate of the magnitude of the rotation rate may be obtained as follows: Assuming some portion of the center-of-mass OAM pPlanck-constant-over-2-$\hbar p$ of the ground state is transferred to the localized state via the weak measurement, we find that $mNR|{ds\over dt}|\leq\hbar|p|$, with ${ds\over dt}$ the velocity of the center-of-mass of the localized state, and the rotation rate obeys $|\Omega|={1\over R}|{ds\over dt}|\leq{\hbar|p|\over mNR^{2}}$. In our scaling the unit for time is $t_{s}={\hbar\over 2mR^{2}}$, so the scaled rotation rate obeys $|v|=|\Omega|t_{s}\leq{2|p|\over N}$, or $|v|\leq 2$ for the current example. Figure 4 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 3(b). In this case the uncertainty is greater than the expected width $w\sim 1$ for a two-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates with a velocity $v\sim 1$ Neq2 , and it persists over several rotations around the ring. The initial position of the centroid of the localized solution is set by the outcome of the initial measurement, $\theta^{(0)}=0$ in Fig. 4(b). In addition, if one selects a value for $\theta_{1}\neq 0$ in Fig. 4 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior.
@@
The above argument may appear counter-intuitive since one would expect a localized initial state to expand and produce interference around the ring, giving rise to complicated dynamics. However, it is possible to construct solutions to the single particle Schrödinger equation for a ring that are localised, non-dispersive, and rotating that have non-zero angular momentum Qin . Adapted to the solution of Eq. (13) these solutions take the form
@@
is a physically allowed single-valued wave function. We have verified numerically that when the weak measurement is performed on the two-particle ground state using Eq. (11), the initialized wave function is well approximated near the origin by the above wave function $\psi(\theta_{1},t=0)$. The physical relevance of this is that, given our choice of measurement distribution in Eq. (12) with $n=1$ and $\theta^{(0)}_{2}=0$, the weak measurement dominantly excites a travelling wave solution of the ring that does not spread. In addition, because of the linearity of the single-particle Schrödinger equation, we can always construct superpositions of states such that we have a non-zero background as appears in the simulations.

### v4

# Simulation of time-crystal-like behavior for a few boson chiral soliton model in a ring

@@ Abstract
We present numerical simulations for a chiral soliton model with $N=2,3$ bosons in a ring, this being a few-particle version of our previous mean-field model for a quantum time crystal OW19 . Following Syrwid, Kosior, and Sacha (SKS), the notion is that a precise position measurement of one particle can lead to spontaneous formation of a bright soliton that in a time crystal should rotate intact for at least a few revolutions around the ring SKS_EPL . In their work SKS find spontaneous formation of a soliton due to the position measurement, but quantum fluctuations cause the soliton to subsequently decay before it has a chance to perform even one revolution of the ring. Based on this they conclude that time crystal dynamics are impossible for Wilczek’s model of a bright soliton in a ring. In contrast, for our few boson chiral soliton model, and allowing for imprecise (weak) measurements of the particle position, we show that time-crystal-like behavior is possible allowing for several revolutions of the spontaneously formed soliton around the ring.
@@ I Introduction
A genuine time crystal involves a Hamiltonian system that exhibits sustained periodic motion even in its lowest-energy state. To the best of our knowledge there is only one example of such a system, that involves long range interactions, but with no physical implementation to date KozKyr19 . The issue undermining Wilczek’s initial proposal for a genuine time crystal was nicely illustrated in a paper by Syrwid, Kosior, and Sacha (SKS) in which they simulated N-bosons in a ring that is pierced by a flux tube, and which was initiated from the N-particle ground state which shows no signatures of rotation or localization SKS_EPL . What they showed was that precise position measurement of one particle could indeed lead to the spontaneous formation of an $(N-1)$-particle soliton that is rotating, and they suggested that if this initial soliton could survive several revolutions around the ring then one could at least speak of time-crystal-like behavior for the Wilczek model. However, SKS found that quantum fluctuations cause this initial soliton to subsequently decay before it has a chance to perform even one revolution of the ring: They presented results for particle numbers in the range $N=10-60$, and the lifetime $t_{c}$ of the soliton was found to increase with $N$. We observed the same behavior for a non-chiral version of our model with $N=3$, both for precise and imprecise position measurements (with an uncertainty in the position measurement), the lifetime being longer for the imprecise measurement. The conclusion of SKS remains intact that Wilczek’s model can realize neither a genuine time crystal nor time-crystal-like behavior SKS_EPL .
@@
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 . This paper was followed by discussion in the literature with SKS SKS_comment -OW_lack , culminating in their paper cited above SKS_EPL . The goal of the present paper is to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time-crystal-like behavior is possible in our model. In particular, we find the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson soliton can persist over several revolutions of the ring. We find that for an imprecise or weak position measurement time-crystal-like behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS SKS_EPL . We therefore argue that our chiral model can display time-crystal-like behavior when weak position measurements are employed.
The chiral model we discuss here has recently been experimentally realised in the meanfield limit by Frölian et al frolian and theoretically studied in edmonds ; valenti ; chisholm . This chiral scenario with its density dependent nonlinear gauge potential opens up a number of intriguing possibilities to study topological gauge theories which can emerge as the low-energy limit of strongly correlated systems such as for instance the Chern-Simons theory of fractional quantum Hall phenomena. The experimental realisation of the chiral nonlinear situation frolian utilises a weakly interacting Bose-Einstein condensate where two internal atomic states are coupled by a laser. The experimental protocol relies on the fact that scattering lengths can be different for the two bare internal states. This in turn results in the corresponding atom-photon dressed states having momentum dependent effective collisional interaction strengths, which can be mapped onto a density dependent gauge potential.
@@
The remainder of this paper is organized as follows: Section II describes the few boson version of our chiral soliton model along with a description of the parameters involved. The case of $N=2$ particles is covered in Sec. III and shows the ideas involved in a transparent manner. Section IV contains the main results and covers the three particle case that demonstrates that our few boson chiral soliton model can display time-crystal-like behavior in which a two-boson soliton state Jackiw can execute several revolutions of the ring, and the soliton lifetime is explored as a function of the position uncertainty of the position measurement. Summary and conclusions are given in Sec. V.
@@ II.1 N-particle Schrödinger equation
$i{\partial\Psi\over\partial t}$ $=$ $\sum_{j=1}^{N}\left(-i{\partial\over\partial\theta_{j}}-A(\theta_{1},\theta_{2},\ldots\theta_{N})\right)^{2}\Psi$ (1)
$+{g\over 2}\rho(\theta_{1},\theta_{2},\ldots\theta_{N})\Psi.$
@@ II.2 Model parameters
In our previous mean-field model the chiral soliton model was derived starting from the spinor Schrödinger equation for a system of two-level bosonic atoms trapped in a ring that are dipole coupled using a laser beam carrying orbital angular momentum (OAM) characterized by the winding number $\ell$. By preparing the atoms in the appropriate dressed state a scalar field analysis may be obtained in which the atoms are subject to a vector potential of the form in Eq. (2) with $A^{(0)}={\ell\over 2}$. Moreover, the density-dependent gauge potential results from a collision-induced detuning, with the consequence that the dimensionless parameter $\kappa\propto\ell$. This means that if the winding number of the laser beam $\ell=0$, then $\kappa=0$ and there will be no chiral dynamics and no possibility for time-crystal-like behavior. The term proportional to the dimensionless parameter $g$ describes energy-shifts due to many-body interactions.
@@ III.1 Two-particle Schrödinger equation
$i{\partial\Psi\over\partial t}$ $=$ $\sum_{j=1}^{2}\left(-i{\partial\over\partial\theta_{j}}-{\ell\over 2}-\kappa\eta(\theta_{1}-\theta_{2})\right)^{2}\Psi$ (6)
$+{g\over 2}\eta(\theta_{1}-\theta_{2})\Psi.$
@@
$i{\partial\Psi\over\partial t}$ $=$ $\bigg{[}-2{\partial^{2}\over\partial x^{2}}-{1\over 2}{\partial^{2}\over\partial s^{2}}+2i\left({\ell\over 2}+\kappa\eta(x)\right){\partial\over\partial s}$ (7)
$+2\left({\ell\over 2}+\kappa\eta(x)\right)^{2}+{g\over 2}\eta(x)\bigg{]}\Psi.$
@@ III.2 Two-particle ground state
In the remainder of this section we choose $g=-2$ as an illustrative example, though similar results could be obtained for different combinations of parameters. Figure 2 shows some more detail of the solution for this choice. In particular, Fig. 2(a) shows the scaled energy versus $p$ and reveals that the ground state occurs for $p=\ell=2$, and (b) shows a line plot of the ground state profile $|\varphi(x)|^{2}$ (properly normalized) versus $x$ for $p=2$. Furthermore, the width (full-width at half-maximum) of the ground state profile in Fig. 2(b) with respect the the relative coordinate $x$ of the two particles is $w\sim 1.6$, the parameter $w$ characterizing the expected width of any localized two-particle solutions that may arise.
Figure 3: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2})|^{2}$, and (b) color coded plot of the probability density $p(\theta_{1},t)$ for a measurement with $n=50$ that localizes particle $2$ with an uncertainty $\Delta\theta\sim 0.34$. This example does not display time-crystal-like behavior. Parameters are the same as as in Fig. 1 along with $g=-2$.
Figure 4: Color coded plot of the probability density $p(\theta_{1},t)$ for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2.3$ that dominantly couples the remaining particle after the measurement of particle 2 to a non-dispersing solution for the ring. This example does display time-crystal-like behaviour. Parameters are the same as as in Fig. 1 along with $g=-2$.
III.3 Time-crystal-like dynamics from measurement
To proceed we first recognize that the ground state represented by the probability density in Fig. 2(b) does not correspond to a localized state on the ring since $|\varphi(x)|^{2}$ depends only on the relative coordinate $x$, and there is no information concerning the COM position $s$ of the particle. On the other hand, since the ground state has a winding number $p\neq 0$ it does represent a persistent flow RyuAndCla07 . We can probe further and reconstruct the ground state two-body wave function $\Psi(\theta_{1},\theta_{2})$ (at $t=0$), where $|\Psi(\theta_{1},\theta_{2})|^{2}$ is shown in Fig. 3(a). From this it is straightforward to see that the probability density of finding the particle at $\theta_{1}$ for any value of $\theta_{2}$,
@@ III.3 Time-crystal-like dynamics from measurement
where the function $G(\theta_{2}-\theta_{2}^{(0)})$ reflects the uncertainty of the position measurement of particle $2$, and is a distribution which obeys the periodicity of the ring, and with an adjustable width $\Delta\theta$ (full-width at half-maximum of $|G(\theta)|^{2}$). Here for illustration we employ the following distribution
@@
with ${\cal N}$ a normalization constant. In the limit $n\gg 1$ this distribution approaches a delta-function which corresponds to the precise measurement model adopted in the paper of SKS, whereas for $n=1$ the measurement uncertainty is around $\Delta\theta\sim 2.3$. After the measurement, precise or imprecise, the wave function $\psi(\theta_{1},t=0)$ (suitably normalized) for the remaining particle is given by Eq. (11), and subsequent time development is governed by the single-particle Schrödinger equation
@@
Taking into account the $e^{ip\theta_{1}/2}$ angular variation associated with the ground state solution (8), the general solution of this equation is
$\psi(\theta_{1},t)={e^{ip\theta_{1}/2}\over\sqrt{2\pi}}\sum_{j=-\infty}^{\infty}c_{j}e^{-ij^{2}t},$ (14)
with expansion coefficients $c_{j}$. This general solution reveals the fact that the single-particle quantum dynamics must be periodic in time with period $2\pi$.
@@
Figure 3(b) shows a simulation of the single-particle probability density $p(\theta_{1},t)=|\psi(\theta_{1},t)|^{2}$ versus $t$ and $\theta_{1}$ according to Eq. (13) with $\theta_{2}^{(0)}=0$ without loss of generality, and $n=50$ so that $\Delta\theta\sim 0.34$. We chose $n=50$ as this approaches a precise measurement as closely as possible within our non-local model with $q=50$ in Eq. (12). In this case the uncertainty is much less than the expected width $w\sim 1.6$ for a two-particle localized state, so the measurement is expected to be very disruptive. We note that, as expected, the probability density varies periodically in time with period $2\pi$. This temporal periodicity might suggest time-crystal-like behavior but this is not the case since the initial probability density profile $p(\theta_{1},t=0)$ becomes rapidly distorted as time increases, not simply translated in time as expected for a time crystal. This is in perfect keeping with the findings of SKS that too precise a position measurement of one particle can produce a localized state initially but this will become rapidly dispersed before the initial localized solution can exhibit even a single rotation around the ring.
@@
For time-crystal-like behavior what is required is that the center-of-mass of the initial localized state translates in time while maintaining its shape for several transits of the ring. An estimate of the magnitude of the rotation rate may be obtained as follows: Assuming some portion of the center-of-mass OAM pPlanck-constant-over-2-$\hbar p$ of the ground state is transferred to the localized state via the weak measurement, we find that $mNR|{ds\over dt}|\leq\hbar|p|$, with ${ds\over dt}$ the velocity of the center-of-mass of the localized state, and the rotation rate obeys $|\Omega|={1\over R}|{ds\over dt}|\leq{\hbar|p|\over mNR^{2}}$. In our scaling the unit for time is $t_{s}={2mR^{2}/\hbar}$, so the scaled rotation rate obeys $|v|=|\Omega|t_{s}\leq{2|p|\over N}$, or $|v|\leq 2$ for the current example. Figure 4 shows an example of time-crystal-like behavior using $n=1$ and $\Delta\theta\sim 2.3$, which means a much weaker measurement than in Figure 3(b). In this case the uncertainty is greater than the expected width $w\sim 1.6$ for a two-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates with a velocity $v\sim 1$ Neq2 , and it persists over several rotations around the ring - an infinite number for the $N=2$ example since solutions of Eq. (13) must be periodic with period $2\pi$. The initial position of the centroid of the localized solution is set by the outcome of the initial measurement, $\theta^{(0)}=0$ in Fig. 4(b), and the centroid simply translates at velocity $v$ as time advances. This example therefore exhibits time-crystal-like behavior.
@@
The above argument may appear counter-intuitive since one would expect a localized initial state to expand and produce interference around the ring, giving rise to complicated dynamics. However, it is possible to construct solutions to the single-particle Schrödinger equation for a ring that are localised, non-dispersive, and rotating that have non-zero angular momentum Qin . Adapted to the solution of Eq. (13) these solutions take the form
@@
is a physically allowed single-valued wave function that represents an excited state. We have verified numerically that when the weak measurement is performed on the two-particle ground state using Eq. (11), the initialized wave function may be approximated near the origin by the above wave function $\psi(\theta_{1},t=0)$. The physical relevance of this is that, given our choice of measurement distribution in Eq. (12) with $n=1$ and $\theta^{(0)}_{2}=0$, the weak measurement can dominantly excite a travelling wave solution of the ring that is initially centered at $\theta_{1}=0$ and that does not spread significantly over several transits of the ring. In addition, because of the linearity of the single-particle Schrödinger equation (13), we can always construct superpositions of states such that we have a non-zero background as appears in the simulations.
III.4 Role of the measurement strength
The results shown in Figs. (3) and (4) show the extremes of strong and weak measurements, respectively, and how they impact the quantum dynamics as reflected in the probability density $p(\theta_{1},t)$ of our system. Next we provide some details of the transition between these extremes and how this may be quantified.
To proceed we first consider the strong measurement example shown in Fig. 3(b), and in Fig. 5(a) we plot the corresponding probability density $p(\theta_{1},t)$ versus $t=[0,2\pi]$ for $\theta_{1}=0$ (solid line) and $\theta_{1}=\pi/2$ (dashed line). Choosing $\theta_{1}=0$ the time variation of the probability density clearly displays periodic behavior with large contrast between the minimum and maximum values, whereas for $\theta_{1}=\pi/2$, while the periodicity is still evident, the contrast between the minimum and maximum values is reduced by a factor of two. This dependence of the temporal variation of the probability density upon the choice of spatial angular position $\theta_{1}$, with concomitant variation in contrast, is a feature that arises when using a strong measurement for which $\Delta\theta/w<1$, that is the angular uncertainty $\Delta\theta$ of the measurement is less than the width $w$ of the $N=2$ solution. The same conclusion arises when considering the spatial profile of the probability density for given times, and Fig. 5(b) shows the probability density versus $\theta_{1}=[-\pi,\pi]$ for $t=0,2\pi$ (solid line), $t=\pi/2$ (dash-dot line), and $t=\pi$ (dash line). The solid line shows both the initial localized probability density created by the measurement (plus the same superposed after $t=2\pi$), whereas at $t=\pi/2$ (dash-dot line) the probability density profile has significantly lower contrast. The spatial profile of the probability density is also the same as the input at $t=\pi$, it is just shifted. In contrast, for the example of a weak measurement with $\Delta\theta/w>1$ as illustrated in Fig. 4, the contrast in the spatial probability density $p(\theta_{1},t)$ for a given time remains constant to within a few percent.
The above discussion highlights that the contrast in the spatial profile of the probability density can serve as a quantitative measure of time-crystal-like behavior in our system. Specifically, following SKS SKS_EPL we here define the time dependent contrast as
$C(t)={max_{\theta_{1}}[p(\theta_{1},t)]-min_{\theta_{1}}[p(\theta_{1},t)]\over max_{\theta_{1}}[p(\theta_{1},t)]+min_{\theta_{1}}[p(\theta_{1},t)]},$ (17)
and we expect time-crystal-like behavior to be absent when $C_{min}=min_{t}[C(t)]$ is quite a bit smaller than $C(0)$, whereas time-like behavior is possible when $C_{min}=min_{t}[C(t)]\simeq C(0)$. Figure 5(c) shows the normalized contrast $C_{min}/C(0)$ versus $n=[1,20]$ from our simulations, and we see that this has dropped to $0.5$ for $n=5$, for which $\Delta\theta/w=0.65$. Consistent with this strong measurement we then find no time-crystal-like behavior for $n=5$ as illustrated in Fig. 5(d) which shows the corresponding color-coded plot of the probability density $p(\theta_{1},t)$ with concomitant degradation in the solution profile as it rotates in comparison to Fig. (4).
As noted earlier, for $N=2$ the quantum dynamics after the measurement must be periodic in time with period $2\pi$, so there is no true decay in this system. On the other hand, following SKS SKS_EPL , we can estimate the lifetime $t_{c}$ of the rotating solution as the minimum time for which the normalized contrast falls below $0.5$. Due to the periodicity any relevant lifetime must be less than $2\pi$. Based on this definition we find from simulations that the lifetime for $n=50$ in Fig. 3(b) is $t_{c}=1$, whereas for $n=5$ in Fig. 5(d) it is $t_{c}=\pi/2$, both of which are shorter than the period $2\pi$ of one rotation. In contrast, for $n=1$ the normalized contrast never falls below $0.9$, so the rotating solution remains largely intact for all times with some small superposed oscillations present, see Fig. 4. It is in this sense that our $N=2$ chiral soliton model can show time-crystal-like behavior to a good approximation.
Figure 5: (a) Probability density $p(\theta_{1},t)$ versus $t=[0,2\pi]$ for $\theta_{1}=0$ (solid line) and $\theta_{1}=\pi/2$ (dashed line); (b) probability density versus $\theta_{1}=[-\pi,\pi]$ for $t=0,2\pi$ (solid line), $t=\pi/2$ (dash-dot line), and $t=\pi$ (dash line); (c) the normalized contrast $C_{min}/C(t=0)$ as a function of $n$; (d) Color coded plot of the probability density for $n=5$. Parameters are the same as as in Fig. 1 along with $g=-2$.